# The Dark Side of LLMs: Agent-based Attacks for Complete Computer Takeover

### v4

@@ Abstract
The rapid adoption of Large Language Model (LLM) agents and multi-agent systems enables remarkable capabilities in natural language processing and generation. However, these systems introduce unprecedented security vulnerabilities that extend beyond traditional content generation attacks to system-level compromise. This paper presents a comprehensive evaluation of the security of LLMs used as reasoning engines within autonomous agents, highlighting how they can be exploited as attack vectors capable of achieving complete computer takeover. We focus on how different attack surfaces and trust boundaries – Direct Prompt Injection, RAG Backdoor, and Inter Agent Trust – can be leveraged to orchestrate such takeovers. We demonstrate that adversaries can effectively coerce popular LLMs (including GPT-4, Claude-4 and Gemini-2.5) into autonomously installing and executing malware on victim machines. Our evaluation of 18 state-of-the-art LLMs reveals an alarming scenario: 94.4% of models succumb to Direct Prompt Injection and 83.3% are vulnerable to the more stealth and evasive RAG Backdoor Attack. Notably, we tested trust boundaries within multi-agent systems, where LLM agents interact and influence each other, and we revealed a critical security flaw: LLMs which successfully resist direct injection or RAG backdoor will execute identical payloads when requested by peer agents. Our findings show that 100.0% of tested LLMs can be compromised through Inter-Agent Trust Exploitation attacks and that every model exhibits context-dependent security behaviors that create exploitable blind spots. Our results also highlight the need to increase awareness and research on the security risks of LLMs, showing a paradigm shift in cybersecurity threats, where AI tools themselves become sophisticated attack vectors.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly accelerated the implementation of artificial intelligence across diverse domains, and the rise of LLM-based agents capable of tackling complex and safety-critical real-world tasks including finance [30], cybersecurity analysis [3], healthcare [1], and autonomous driving [19].
@@
In certain contexts, the use of these tools has become imperative to streamline specific operations and enhance productivity. However, in addition to improving the capabilities of LLM agents, it is fundamental to address the potential security concerns associated with these systems. For example, shopping agents can search for, monitor, and notify users about deals on requested products. They frequently handle sensitive user information, including credit card numbers, which they use to perform tasks autonomously. The disclosure of private information of the customer by the agent, while completing the autonomous web shopping, would result in severe damage.
Moreover, to solve particular and non-trivial tasks, the agentic pipeline is often supported by retrieving knowledge from a Retrieval-Augmented Generation (RAG) [15] knowledge base, a state-of-the-art technique designed to mitigate LLM limitations such as outdated knowledge, hallucinations, and domain-specific gaps. An agentic RAG [26] based on the ReAct paradigm [35] usually operates through several key steps when solving a task: (i) defining roles and behaviors via a system prompt; (ii) receiving user instructions and task details; (iii) retrieving relevant information from an external database; (iv) planning actions based on the retrieved information and the prior context; (v) executing actions using external tools.
@@
While each of these steps enables the agent to perform highly complex tasks, they also provide adversaries with multiple new attack surfaces to compromise the agent or, even more dangerously, to gain full control over the agent host platform. Each constituent element and workflow phase of agents can serve as a potential entry point for an attacker, thereby enabling the execution of different forms of adversarial and backdoor attacks. Furthermore, the transition from isolated LLM agents to modern multi-agent systems introduces novel techniques and trust boundaries for the exploitation of impersonation, task tampering, and unauthorized privilege escalation threats.
@@
In this work, we aim to evaluate the intrinsic security mechanisms of LLMs, specifically their ability to detect and resist textual instructions that violate cybersecurity norms. This analysis is no longer merely theoretical nor limited to traditional prompt-based interactions because LLMs are increasingly used not just to generate natural language responses but also to act as reasoning engines for autonomous agents. As such, any failure of an LLM to recognize and reject malicious instructions can have real-world consequences, elevating the security of LLMs behavior from a language modeling concern to a critical system safety issue. More in detail, we show that different attack surfaces and trust boundaries within LLM agents can be abused to deceive the LLM and trigger the execution of harmful code, potentially gaining control over the agent’s hosting platform (hereafter referred to as the victim machine). This process, depending on the agent structure and attack technique, often occurs without the knowledge or awareness of the end user, who ultimately becomes a victim of the attack.
@@
Furthermore, we present a pivotal result related to trustiness in multi-agent systems. We observed that, in instances where some LLMs (see Section 4 for further details) are capable of identifying and rejecting malicious classified commands – retrieved from any visible or hidden step of the workflow – these same models will execute those precise commands if they are propagated by another agent within a multi-agent system. In this scenario, the LLM treats the input as trustworthy because it originates from a peer entity.
@@
These discoveries highlight a significant shift in the cybersecurity landscape: cyberattack frontiers are moving away from traditional techniques, such as phishing, infected USB devices, or direct exploitation of operating system vulnerabilities, toward novel attack vectors that leverage commonly used AI tools and multi-agent systems. These attacks also imply a serious threat to users because AI-based tools are typically designed to be highly accessible and user-friendly, requiring minimal to no technical expertise. This significantly lowers the barrier to conducting sophisticated attacks, expanding the attack surface and allowing even low-skilled adversaries to engage in malicious behavior.
@@
  • We present the first systematic study on the feasibility of using LLM-powered systems as an attack vector. We demonstrate how LLM can be exploited to achieve complete computer takeover, moving beyond content generation attacks to system-level compromise. Our evaluation spans 18 state-of-the-art LLMs across three distinct attack surfaces and the corresponding trust boundaries: Direct Prompt Injection, RAG Backdoor, Inter-Agent Trust.
  • We show how adversaries can compromise the agent knowledge bases provided through RAG and trigger malicious behavior during routine agent operations that can affect system and user security and privacy while pursuing intended tasks. Our RAG Backdoor Attack successfully compromise 83.3% of the tested models.
  • We reveal a critical vulnerability in multi-agent systems where LLMs treat peer agents as inherently trustworthy, bypassing safety mechanisms designed for human-AI interactions. Our findings show that 100.0% of the tested models execute malicious commands when requested by peer agents, even when they successfully resist identical direct or indirect command injection.
  • Our analysis demonstrates that LLM-based attacks require minimal technical expertise while achieving maximum impact through the deployment of autonomous malware.
The remainder of the paper is organized as follows. Section 2 provides the necessary background on agentic AI systems and the technical foundations relevant to our work. Section 3 details the methodology adopted for our analysis, including the threat modeling process and the rationale behind key design decisions. Section 4 presents our experimental findings and discusses the observed vulnerabilities, while Section 5 shows a comprehensive analysis of the sensitivity of each model as malicious prompt changes. Section 6 addresses the broader ethical implications of this research, particularly the real-world risks associated with the discovered vulnerabilities. In Section 7 we review related work and contextualize our contributions within the existing literature. Finally, we draw our conclusions in Section 8.
@@ 2.1 Agentic AI systems and LLM Agents
An agent [29] is defined as a computer system situated in an environment that is capable of acting autonomously in its context to achieve its delegated objectives. Autonomy means the ability and requirements to decide how to act to achieve a goal. An agent that can perceive its environment, react to changes that occur in it, take the initiative, and interact with other systems (like other agents or humans) is called an intelligent agent or Agentic AI. Effective memory management improves an agent’s ability to maintain context, learn from past experiences, and make more informed decisions over time. In recent developments, Agentic AI systems are evolving from isolated, task-specific models into dynamic and multi-agent ecosystems (MAS).
As pointed out in [31], the growth of LLMs has culminated in the emergence of LLM agents. They use LLMs as reasoning and planning cores to decide the control flow of an application while maintaining the characteristics of traditional intelligent agents. LLM agents can invoke external tools for the resolution of specific tasks and can decide whether the generated answer is sufficient or if further work is necessary.
Figure 1: Intelligent Agent Structure [31]
An emerging class of LLM agents is agentic RAG, which employs the RAG paradigm [15] to reduce hallucinations and improve the domain-specific expertise of an LLM.
2.2 Attacks to the LLMs
@@
When an attacker interacts directly with the chatbot and embeds malicious instructions in the dialogue, the attack is referred to as a direct prompt injection. In contrast, an indirect prompt injection occurs when the attacker manipulates external content, such as documents or data sources, that the AI system later processes, causing it to behave in an unintended way [8].
LLM Backdoor Attacks. These attacks aim to inject a backdoor into a model, causing it to behave normally on benign inputs but produce malicious outputs when triggered by a specific pattern or rule. The goal of traditional backdoor attacks is to build shortcuts between trigger and target labels in specific downstream tasks for language models [9, 12, 17]. There are two commonly used techniques for injecting backdoors: data poisoning and weight poisoning.
@@
LLM Agent Backdoor Attacks. Backdoor attacks on LLM agents, also referred to as indirect prompt injection attacks, differ from those targeting traditional LLMs, as agents perform multi-step reasoning and interact with the environment to acquire external information before generating the output [10]. As pointed out in [34], more opportunities for sophisticated attacks, such as query-attack, observation-attack, and thought-attack, are created by this extended workflow of LLM agents. In fact, these attacks can be carried out on any hidden step in the reasoning, planning, and action of the agents without compromising the final output and remaining stealthy for the user who became an unintentional victim.
@@ 3 Exploiting LLM Agent-based Attack
Our goal is to demonstrate that intelligent systems introduce novel and various attack surfaces and corresponding trust boundaries within LLM that can be abused by a malicious actor to transform these tools into a modern attack vector. In our designed scenario, any failure of an LLM to recognize and reject malicious instructions implies that the adversary is able to gain full control over the agent host platform by coercing the model into installing and executing malware.
@@ 3.1 Threat Model
Black-box setting of the agent systems. We assume a black-box setting where attackers do not have access to: (i) the internal parameters and weights of the underlying LLMs, (ii) the RAG embeddings model, (iii) and the employed retrieval techniques.
@@
Assumption for the attacker capabilities. We strictly follow the standard threat model assumptions for RAG Backdoor attacks. We assume that the attacker has partial access to the RAG database, which means they only have the capability to inject some malicious text into the external source to create a poisoned database [36, 24, 6, 5]. This assumption aligns with practical scenarios where the agent’s external knowledge source unit is hosted by a third-party retrieval service or directly leverages an unverified knowledge base.
@@
Attacker goal. The attacker pursues two adversarial goals. The attacker’s primary goal is to misdirect the agent to execute specific actions that align with the attacker’s intent but are unintended by the user. The second goal is to maintain the perceived integrity of the output: whether or not the malicious actions are executed, the user should consistently receive the expected response, with no visible signs of compromise.
In our tests, the attacker’s intent is to ensure that malware is successfully installed on the victim’s machine whenever the agent retrieves and processes the malicious command at any point in its workflow.
Agent architecture assumption. To support our analysis, we developed agentic RAG and LLM agents with terminal access. As reported in many recent works [25, 7, 27], websites [11] and GitHub projects [13, 20, 2], giving agents access to a bash environment or system shell has become a common practice to improve their autonomy.
3.2 Agent and Adversarial Payload Design
In our analyses, we want to test both: (i) different attack techniques in diverse categories of modern AI agents (ii) the sensitivity of each LLM to such attacks.
@@ 3.2 Agent and Adversarial Payload Design
To achieve our goal, we developed the necessary agents using state-of-the-art framework for the creation of application powered by LLM: LangChain and LangGraph [4, 21]. The relevant agent tool implemented are: (i) a retrieval tool which is in charge of searching for relevant information in the RAG knowledge base; (ii) a tool that allows the agent to interact with a system terminal; (iii) in the context of Agentic AI systems, we also implemented a tool to allow agents to communicate with each other.
@@
The malware to be installed, based on Meterpreter [22], initiates an outbound TCP connection to the attacker’s machine and enables remote access to the victim’s machine. The resulting reverse shell not only provides a wide range of commands to perform post-exploitation operations, but is also executed entirely in memory. This in-memory execution avoids writing files to disk, thereby significantly reducing the likelihood of detection by conventional security mechanisms.
@@
The full malicious prompt sent to the agent consists of three parts: (i) a payload containing the Base64-encoding of the Meterpreter malware; (ii) a sequence of instructions that prompt the agent to decode the payload and execute it in the background mode, namely, a command pipe; (iii) a message that contains one or more sentences designed to “fool” the agent to execute the command pipe while completing the original user task. To evaluate the system’s response to different prompts, we designed three command pipes and two malicious messages, while maintaining the same underlying payload (see Section 5.1 for more details about command pipes and messages). The malicious prompt is then delivered to the agents in several ways depending on the specific attack technique (see Section 4 for more details).
3.3 Synthetic applications overview

### v5

@@ Abstract
The rapid adoption of Large Language Model (LLM) agents and multi-agent systems enables remarkable capabilities in natural language processing and generation. However, these systems introduce security vulnerabilities that extend beyond traditional content generation to system-level compromises. This paper presents a comprehensive evaluation of the LLMs security used as reasoning engines within autonomous agents, highlighting how they can be exploited as attack vectors capable of achieving computer takeovers. We focus on how different attack surfaces and trust boundaries can be leveraged to orchestrate such takeovers. We demonstrate that adversaries can effectively coerce popular LLMs into autonomously installing and executing malware on victim machines. Our evaluation of 18 state-of-the-art LLMs reveals an alarming scenario: 94.4 % of models succumb to Direct Prompt Injection, and 83.3 % are vulnerable to the more stealthy and evasive RAG Backdoor Attack. Notably, we tested trust boundaries within multi-agent systems, where LLM agents interact and influence each other, and we revealed that LLMs which successfully resist direct injection or RAG backdoor attacks will execute identical payloads when requested by peer agents. We found that 100.0 % of tested LLMs can be compromised through Inter-Agent Trust Exploitation attacks, and that every model exhibits context-dependent security behaviors that create exploitable blind spots.
keywords:
[…]
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly accelerated the implementation of artificial intelligence across diverse domains, including the rise of LLM-based agents capable of tackling complex and safety-critical real-world tasks, such as finance [35], cybersecurity analysis [5], healthcare [1], and autonomous driving [21].
@@
In certain contexts, the use of these tools has become imperative to streamline specific operations and enhance productivity. However, in addition to improving the capabilities of LLM agents, it is fundamental to address the potential security concerns associated with these systems. For example, shopping agents can search for, monitor, and notify users about deals on requested products. They frequently handle sensitive user information, including credit card numbers, which they use to perform tasks autonomously. The disclosure of private information of the customer by the agent, while completing autonomous web shopping, would result in severe damage.
@@
Moreover, to solve particular and non-trivial tasks, the agentic pipeline is often supported by retrieving knowledge from a Retrieval-Augmented Generation (RAG) [17] knowledge base, a state-of-the-art technique designed to mitigate LLM limitations such as outdated knowledge, hallucinations, and domain-specific gaps. An agentic RAG [31] based on the ReAct paradigm [40] usually operates through several key steps when solving a task: (i) defining roles and behaviors via a system prompt; (ii) receiving user instructions and task details; (iii) retrieving relevant information from an external database; (iv) planning actions based on the retrieved information and the prior context; (v) executing actions using external tools. While each of these steps enables the agent to perform highly complex tasks, they also provide adversaries with multiple new attack surfaces to compromise the agent or, even more dangerously, to compromise or gain full control over the agent host platform. Each constituent element and the workflow phases of agents can serve as potential entry points for an attacker, thereby enabling the execution of different forms of adversarial and backdoor attacks. Furthermore, the transition from isolated LLM agents to modern multi-agent systems introduces novel techniques and trust boundaries for the exploitation of impersonation, task tampering, and unauthorized privilege escalation threats.
@@
In this work, we evaluate the intrinsic security mechanisms of LLMs, specifically their ability to detect and resist textual instructions that violate cybersecurity norms. This analysis is no longer merely theoretical, nor is it limited to traditional prompt-based interactions because LLMs are increasingly used not just to generate natural language responses but also to act as reasoning engines for autonomous agents. As such, any failure of an LLM to recognize and reject malicious instructions can have real-world consequences, elevating the security of LLMs behavior from a language modeling concern to a critical system safety issue. More in detail, we analyze different attack surfaces and trust boundaries within LLM agents and find exploitable spots that can be abused to deceive the LLM and trigger the execution of harmful code, potentially gaining control over the agent’s hosting platform (hereafter referred to as the victim machine). We also analyze different agents structure and attack techniques that an attacker can exploit without the knowledge or awareness of the end user, who ultimately becomes a victim of the attack.
@@
Furthermore, we present a pivotal result related to trustiness in multi-agent systems. We observed that, in instances where some LLMs (see Section 4 for further details) are capable of identifying and rejecting malicious commands – retrieved from any visible or hidden step of the workflow – these same models will execute those precise commands if they are propagated by another agent within a multi-agent system.
@@
These discoveries highlight a significant shift in the cybersecurity landscape: cyberattack frontiers are moving away from traditional techniques, such as phishing, infected USB devices, or direct exploitation of operating system vulnerabilities, toward novel attack vectors that leverage commonly used AI tools. These attacks also imply a serious threat to users because AI-based tools are typically designed to be highly accessible and user-friendly, requiring minimal to no technical expertise. This significantly lowers the barrier to conducting sophisticated attacks, expanding the attack surface and allowing even low-skilled adversaries to engage in malicious behavior.
@@
  • We present a systematic study on the feasibility of using LLM-powered systems as an attack vector, demonstrating how an LLM can be exploited to achieve complete computer takeover, moving beyond content generation attacks to system-level compromise. Our evaluation spans 18 state-of-the-art LLMs [^1] At the time of writing, July 2025. across three distinct attack surfaces and the corresponding trust boundaries: Direct Prompt Injection, RAG Backdoor, Inter-Agent Trust Exploitation.
  • We show how adversaries compromising the agent knowledge bases provided through RAG can trigger malicious behavior during routine agent operations, which can affect the security and privacy of the system and users while pursuing intended tasks.
  • We reveal a critical vulnerability in multi-agent systems where LLMs treat peer agents as inherently trustworthy, bypassing safety mechanisms designed for human-AI interactions. Our findings show that 100.0 % of the tested models execute malicious commands when requested by peer agents, even when they successfully resist identical direct or indirect command injections.
  • We conduct a sensitivity analysis to assess how susceptible LLMs are to variations in the prompts formulation. Moreover, we demonstrates that LLM-based attacks for system level compromise require minimal technical expertise while achieving maximum impact through the deployment of autonomous malware.
To systematically evaluate these security concerns and provide empirical evidence for our claims, we structured our investigation around the following research questions:
  • RQ1: LLM Security Mechanisms and Attack Surfaces in Agentic Contexts. Are the intrinsic security mechanisms of LLMs – developed through substantial investment in safety training – sufficient in LLM-based agentic systems with autonomous tool execution capabilities? Can different attack surfaces and trust boundaries be exploited to achieve a complete system-level compromise?
  • RQ2: Unwitting User Victimization. Can users become unwitting victims of attacks executed through LLM-based agentic systems without any direct interaction between the attacker and the agent or the user, in both single-agent and multi-agent contexts?
  • RQ3: Trust Boundary Vulnerabilities in Multi-Agent Systems. Do multi-agent systems exhibit specific security vulnerabilities related to inter-agent trust, and how do these differ from direct attack scenarios?
The remainder of the paper is organized as follows. Section 2 provides the necessary background on agentic AI systems and the technical foundations relevant to our work. Section 3 details the methodology adopted for our analysis, including the threat modeling process and the rationale behind key design decisions. Section 4 presents our experimental findings and discusses the observed vulnerabilities, while Section 5 provide a comprehensive analysis of the sensitivity of each model as malicious prompt changes. Section 6 addresses the impact of this research with a focus on the real-world risks associated with the discovered vulnerabilities. In Section 7 we review related work and contextualize our contributions within the existing literature. Finally, we draw conclusions in Section 8.
@@ 2.1 Agentic AI systems and LLM Agents
An agent [34] is defined as a computer system situated in an environment that is capable of acting autonomously in its context to achieve its delegated objectives. Autonomy means the ability and requirements to decide how to act to achieve a goal. An agent that can perceive its environment, react to changes that occur in it, take the initiative, and interact with other systems (like other agents or humans) is called an intelligent agent or Agentic AI system. Effective memory management improves an agent’s ability to maintain context, learn from past experiences, and make more informed decisions over time. In recent developments, Agentic AI systems are evolving from isolated, task-specific models into dynamic and multi-agent ecosystems (MAS).
As pointed out in [36], the growth of LLMs has culminated in the emergence of LLM agents. They use LLMs as reasoning and planning cores to decide the control flow of an application while maintaining the characteristics of traditional intelligent agents. LLM agents can invoke external tools for the resolution of specific tasks and can decide whether the generated answer is sufficient or if further work is necessary. An emerging class of LLM agents is agentic RAG, which employs the RAG paradigm [17] to reduce hallucinations and improve the domain-specific expertise of an LLM.
Figure 1: (a) generic structure of an intelligent agent and its autonomous interactions with both the environment and external tools [36]. (b) principal attack surfaces affecting LLM agents, such as direct prompt injection and RAG-based knowledge base poisoning, with an example exploit of the agent’s workflow [41].
2.2 Attacks to the LLMs and LLM Agents
@@
When an attacker interacts directly with the chatbot and embeds malicious instructions in the dialog, the attack is referred to as a direct prompt injection. In contrast, an indirect prompt injection occurs when the attacker manipulates external content, such as documents or data sources, that the AI system later processes, causing it to behave in an unintended way [12].
LLM Backdoor Attacks. These attacks aim to inject a backdoor into a model, causing it to behave normally on benign inputs but to produce malicious outputs when triggered by a specific pattern or rule. The goal of traditional backdoor attacks is to build shortcuts between trigger and target labels in specific downstream tasks for language models [13, 15, 19]. There are two commonly used techniques for injecting backdoors: data poisoning and weight poisoning.
@@
LLM Agent Backdoor Attacks. Backdoor attacks on LLM agents, also referred to as indirect prompt injection attacks, differ from those targeting traditional LLMs, as agents perform multi-step reasoning and interact with the environment to acquire external information before generating the output [14]. As pointed out in [39], more opportunities for sophisticated attacks, such as query-attack, observation-attack, and thought-attack, are created by this extended workflow of LLM agents. In fact, these attacks can be carried out on any hidden step in the reasoning, planning, and action of the agents without compromising the final output and remaining stealthy for the user who became an unintentional victim. Figure 1 illustrates the architecture of modern intelligent agents and highlights the main attack surfaces introduced by integrating LLMs as decision and reasoning cores.
@@ 3 Exploiting LLM Agent-based Attack
Our goal is to demonstrate that intelligent systems powered by LLM introduce novel and various attack surfaces and corresponding trust boundaries that can be abused by a malicious actor to transform these tools into a modern attack vector. In our designed scenario, any failure of an LLM to recognize and reject malicious instructions implies that the adversary is able to gain full control over the agent host platform by coercing the model into installing and executing malware.
@@ 3.1 Threat Model
Architecture and Black-box settings of agent systems. We assume a black-box setting where attackers do not have access to any of these agent components: (i) the internal parameters and weights of the LLMs, (ii) the RAG embeddings model, (iii) and the employed retrieval techniques.
Moreover, as reported in many recent works [30, 10, 32], giving agents access to a bash environment or system shell has become a common practice to improve their autonomy. Our evaluation employs agents with full bash terminal access, a configuration already adopted by several production systems, such as highly starred GitHub projects [26, 24, 2, 23, 9, 22]. However, the attack vectors demonstrated in this work extend to environments with more restrictive guardrails.
In the context of multi-agent systems, the threat model must be adapted to reflect the architecture and interaction flows between agents and users. While the security assumptions established above for single agents still apply, in the the multi-agent system only one agent interacts directly with the user and the environment (e.g., is granted access to external sources of information). Consequently, this interface agent represents the sole entry point for both legitimate input and potential compromise, whether by direct or indirect means. The remaining agents within the system operate exclusively on information propagated from the interface agent, never interacting with the user.
@@
Assumption of the attacker capabilities. We strictly follow the standard threat model. We assume that the attacker has partial access to the RAG database, which means they only have the capability to inject some malicious text into the external source to create a poisoned database [43, 29, 8, 7]. This assumption aligns with practical scenarios where the agent’s external knowledge source unit is hosted by a third-party retrieval service or directly leverages an unverified knowledge base.
@@
Attacker goal. The attacker pursues two adversarial goals. The attacker’s primary goal is to misdirect the agent to execute specific actions that align with the attacker’s intent but are unintended by the user. The second goal is to maintain the perceived integrity of the output: whether or not malicious actions are executed, the user should consistently receive the expected response, with no visible signs of compromise. In our tests, the attacker’s intent is to ensure that malware is successfully installed on the victim’s machine whenever the agent retrieves and processes the malicious command at any point in its workflow.
3.2 Agents and Adversarial Payloads Design
@@
Our analyses test both: (i) different attack techniques in diverse categories of modern AI agents (ii) the sensitivity of each LLM to such attacks. We developed the necessary agents using the state-of-the-art framework for the creation of applications powered by LLM: LangChain and LangGraph [6, 25]. The relevant agent tool implemented are: (i) a retrieval tool which is in charge of searching for relevant information in the RAG knowledge base; (ii) a tool that allows the agent to interact with a system terminal; (iii) in the context of multi-agent systems, we also implemented a tool to allow agents to communicate with each other.
@@
The final attacker intent is to deploy malware on victim machine. The malware to be installed, based on Meterpreter [27], initiates an outbound TCP connection to the attacker’s machine, which acts as a Command&Control (C&C) server and enables remote access to the victim’s machine. The resulting reverse shell not only provides a wide range of commands to perform post-exploitation operations, but it is also executed entirely in memory. This in-memory execution avoids writing files to disk, thereby significantly reducing the likelihood of detection by conventional security mechanisms.
@@
The malicious prompt sent to the agent consists of three parts: (i) a payload containing the Base64-encoding of the Meterpreter malware; (ii) a sequence of instructions that prompt the agent to decode the payload and execute it in the background mode, namely, a command pipe; (iii) a message that contains one or more sentences designed to “fool” the agent to execute the command pipe while completing the original user task. To evaluate the system’s response to different prompts, we designed three command pipes and two malicious messages, while maintaining the same underlying payload (see Section 5.1 for more details about command pipes and messages). The malicious prompt is then delivered to the agents in several ways depending on the specific attack technique (see Section 4 for more details).
Notably, our prompts are deliberately designed to be model-agnostic, functioning across different LLM architectures without requiring model-specific customization. They exploit reasoning patterns common to LLMs rather than relying on model-specific architectural peculiarities. These prompts demonstrate that the attack surface is not limited to a specific LLM but represents a systemic concern across the entire class of agent-based systems, regardless of the underlying language model architecture or training methodology.
3.3 Synthetic Application and Attack Description